OpenStudy (anonymous):

Can someone help walk me through how to find the derivative of tan(x) using the limit definition?

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0} \frac{ (tanx+h)-\tan(x) }{ h }\]

OpenStudy (helder_edwin):

go ahead. \[ \large (\tan x)'=\lim_{h\to0}\frac{\tan(x+h)-\tan x}{h} \]

OpenStudy (helder_edwin):

check your parenthesis.

OpenStudy (anonymous):

\[\lim_{h \rightarrow }(\frac{ tanx+\tanh }{ 1-tanxtanh }-tanx)/h\]

OpenStudy (anonymous):

oh woops, that's what I meant.

OpenStudy (anonymous):

but then I have no idea how to simplify that.

OpenStudy (helder_edwin):

\[ \large \tan(x+h)=? \]

OpenStudy (anonymous):

recall a trig identity for tan (x+h)...

OpenStudy (anonymous):

\[\frac{ tanx+\tanh }{ 1-tanxtanh}\]

OpenStudy (helder_edwin):

great

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

that's where I'm stuck

OpenStudy (helder_edwin):

so we have \[ \large (\tan x)'=\lim_{h\to0}\frac{\frac{\tan x+\tan h}{1-\tan x\tan h}-\tan x}{h} \]

OpenStudy (anonymous):

yep, that's where I got

OpenStudy (anonymous):

I think I maybe got it

OpenStudy (helder_edwin):

add the fractions in the numerator.

OpenStudy (anonymous):

convert than tans to sin/cos then factor out the sinx/cosx to get all the trig h's in the limit...

OpenStudy (anonymous):

nevermind, that didn't work

OpenStudy (anonymous):

ok

OpenStudy (helder_edwin):

so \[ \large (\tan x)'=\frac{\tan x+\tan h-\tan x+\tan^2x\tan h}{h(1-\tan x\tan h)} \]

OpenStudy (helder_edwin):

right?

OpenStudy (anonymous):

yes.

OpenStudy (helder_edwin):

simplify and factor.

OpenStudy (anonymous):

what am I factoring out? tanh?

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0}\frac{ \tanh(1-\tan ^{2}x) }{ (1_tanxtanh)(h) }\]

OpenStudy (anonymous):

sorry, the denominator should be 1-tanxtanh

OpenStudy (anonymous):

times h

OpenStudy (helder_edwin):

no

OpenStudy (anonymous):

now I'm lost again lol

OpenStudy (helder_edwin):

\[ \large (\tan x)=\lim_{h\to0}\frac{\tan h(1+\tan^2x)}{h(1-\tan x\tan h)} \]

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0}\frac{ (\tanh)(1-\tan ^{2}x) }{ (h)(1-tanxtanh) }\]

OpenStudy (anonymous):

that's what I meant, sorry. I have trouble with the equation maker

OpenStudy (anonymous):

the subtraction sign in the numerator should be an addition @nasryn

OpenStudy (anonymous):

oh.

OpenStudy (helder_edwin):

i know. it is better to use the keyboard directly (if u know latex).

OpenStudy (helder_edwin):

now. let's look at the pieces: \[ \large \lim_{h\to0}\frac{\tan h}{h}=? \]

OpenStudy (anonymous):

the only thing I can think of to do next is to make tanh = sinh/cosh, then cancel the sinh with sinh/h which the liimit of that would be 1

OpenStudy (helder_edwin):

yes.

OpenStudy (anonymous):

I don't know that identity for tanh/h directly

OpenStudy (helder_edwin):

\[ \large \lim_{h\to0}\frac{\tan h}{h}=\lim_{h\to0}\frac{\sin h}{h}\cos h=1\cdot 1=1 \] there is no identity for this limit.

OpenStudy (anonymous):

oh! wow! I feel silly

OpenStudy (helder_edwin):

finally \[ \large \lim_{h\to0}\frac{1}{1-\tan x\tan h}=? \]

OpenStudy (anonymous):

wait.

OpenStudy (anonymous):

would tanh/h be (sinh)/(h)(cosh)

OpenStudy (helder_edwin):

yes. that's what i wrote. \[ \large \frac{\tan h}{h}=\frac{\sin h}{\cos h}\frac{1}{h} \]

OpenStudy (anonymous):

oh ok

OpenStudy (helder_edwin):

sorry. i had a typo.

OpenStudy (anonymous):

so the answer is just 1+tan^2x?

OpenStudy (anonymous):

ohh and that's sec^2(x)!

OpenStudy (helder_edwin):

correct.

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

:) thanks!

OpenStudy (anonymous):

but now how do I get rid of tanxtanh in the denominator?

OpenStudy (anonymous):

oh, just plug in. :P

OpenStudy (helder_edwin):

\[ \large \lim_{h\to0}\frac{1}{1-\tan x\tan h}=\frac{1}{1-\tan x\tan0}=1 \]

OpenStudy (anonymous):

that was rough, I hope that's not on my exam. :P

OpenStudy (helder_edwin):

so \[ \large \lim_{h\to0}\frac{\tan h(1+\tan^2x)}{h(1-\tan x\tan h)}=1+\tan^2x=\sec^2x \]

OpenStudy (helder_edwin):

the best of luck in your test.

OpenStudy (anonymous):

thanks.